OpenStudy (anonymous):

from a base evaluation of 9200 ft, a mountain peak in Colorado rises to a summit elevation of 14,363 ft over a horizontal distance of 15, 841 ft . find the grade of the peak

OpenStudy (anonymous):

round to the nearest tenth as needed

OpenStudy (saifoo.khan):

is the question complete?

OpenStudy (anonymous):

yes

OpenStudy (perl):

use coordinates (0,9200) ( 15841, 14363)

OpenStudy (perl):

[ 14363 - 9200] / ( 15841 - 0

OpenStudy (perl):

= .3259, which is 32.6% grade roughly

OpenStudy (agreene):

\[\frac{\Delta y}{\Delta x}\] \[\frac{14636-9200}{15841-0}=\frac{5436}{15841}=0.3431\]

OpenStudy (perl):

agreene you have a typo

OpenStudy (perl):

its 14363 ,

OpenStudy (saifoo.khan):

what does that mean by final "grade" ?

OpenStudy (agreene):

well blah! stupid typos... they make me sadface

OpenStudy (perl):

i think grade is a steepness , multiply your slope by 100 ,

OpenStudy (saifoo.khan):

the percentage of slope?

OpenStudy (agreene):

and it says "find the grade" if you prefer, the proper mathematical terminology is "Find the Gradient" its the same as slope.

OpenStudy (saifoo.khan):

ohh!!

OpenStudy (perl):

im guessing it means percent

OpenStudy (agreene):

Well, it's the same as slope in 2-space

OpenStudy (saifoo.khan):

then why u are multiply it by 100?

OpenStudy (perl):

because it looks nicer ;)

OpenStudy (perl):

it gives it oomph

OpenStudy (saifoo.khan):

lol, okay.

OpenStudy (agreene):

more info on getting out of simple maths and dealing with Gradients: http://mathworld.wolfram.com/Gradient.html

OpenStudy (perl):

saifoo, because kids dont like decimals. and truckers

OpenStudy (perl):

, because truckers dont want to be told their slope is .32, so we say 32% to make them happy

OpenStudy (saifoo.khan):

lol.

OpenStudy (perl):

so a 45% grade is pretty steep ,

OpenStudy (perl):

err, i mean 100 % grade

OpenStudy (perl):

100 % grade is 45 degrees or pi/4